OpenStudy (anonymous):

solve: (2x+3y=7 (x+y=3

OpenStudy (anonymous):

2x + 3y = 17......... eq1 3x - 5y = -3 .......... eq2 6x + 9y = 51 ........ eq3 = eq1 x 3 -6x + 10y = 6 ....... eq4 = eq2 x -2 19y = 57 .............. eq3 + eq4 y = 3 x = [ 17 - 3(3) ] / 2 . eq1 x = 4 Answer: x=4, y=3